OpenStudy (anonymous):

A bus leaves a terminal traveling at a rate of 45 miles per hour. When the bus has traveled 90 miles, a car leaves the same terminal, traveling at a rate of 60 miles per hour. Which pair of equations model the distances d covered by each vehicle after t hours?

OpenStudy (anonymous):

Bus: d1 = 60t; Car: d2 = 45t – 90

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

yes it is right

OpenStudy (anonymous):

u sure?

OpenStudy (anonymous):

Bus: d1 = 60t + 90; Car: d2 = 45t

OpenStudy (anonymous):

A. Bus: d1 = 45t + 90; Car: d2 = 60t B. Bus: d1 = 60t; Car: d2 = 45t – 90 C. Bus: d1 = 60t + 90; Car: d2 = 45t D. Bus: d1 = 45t + 90; Car: d2 = 60t + 90

OpenStudy (anonymous):

@begzat

OpenStudy (anonymous):

i think its A now

OpenStudy (anonymous):

r*t = d The following equation would be used to find t, the elapsed time for the car to intercept the bus.\[60 t=45t+90\]I hope that helps.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Thank you for the medal.